# Splittings in a complete local ring and decomposition its group of units

### v1

@@ Abstract.
Let $(R,M,k)$ be a complete local ring (not necessarily Noetherian). Then we prove that the natural surjective ring map $R\rightarrow k$ admits a splitting if and only if $\operatorname{char}(R)=\operatorname{char}(k)$. In addition, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then we prove that the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting. As an application of this theorem, we show that the following short exact sequence of Abelian groups:
@@
is always split. We also show with an example that the above exact sequence does not split for most incomplete local rings. In an attempt to prove the above main theorem, we first obtained a characterization for separability which states that an algebraic extension of fields is separable if and only if it is formally étale. In particular, in the characteristic zero, every algebraic extension of fields is formally étale. Next, we show that every field extension over a perfect field is formally smooth.
@@
complete local ring; the group of units; splitting; characteristic methods; (separating) transcendence basis; formally étale ring map; formally smooth ring map; Ext groups; pullbacks in the category of (commutative) rings
@@ 1. Introduction
In mathematics, it is common for a particular case of a hidden fact to appear first of all. Then, as the obstacles and covers are gradually removed, the general form of that phenomenon becomes fully revealed. The same happened in this research work. More precisely, by elementary methods we first showed that for any “finite” local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@
is split. Next, this observation naturally led us to the question of whether the same fact holds for arbitrary (infinite) local rings. Then, after exploring several stages of the level of generalization, first for Artinian local rings, then for local rings whose Jacobson radical are nilpotent, finally as one of the main results of this investigation, we succeeded to show that the same splitting holds for any complete local ring. However, the proof of this general case is technical, and uses methods of commutative algebra, the theory of fields and homological algebra. In fact, we prove a more general result:
@@
In proving this theorem, as is clear from its formulation, the characteristic of the residue field $k$ plays an important role and hence it is necessary to examine it carefully. We know that in any local ring $(R,M,k)$, if $\operatorname{char}(R)=\operatorname{char}(k)$ then there are two cases: $\operatorname{char}(k)=0$ or $\operatorname{char}(k)=p$ for some prime number $p$. However, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the characteristic of $k$ is always a prime number $p$. To prove the above theorem, in the first case, i.e., when $\operatorname{char}(R)=\operatorname{char}(k)$, one of the main ingredients that applied is the following key lemma:
@@
We then obtained the following groundbreaking result, which paves the way for proving Lemma 1.2, especially in the positive characteristic case.
Theorem 1.4.
@@
To prove Theorem 1.1, in the second case, i.e., when $\operatorname{char}(R)\neq\operatorname{char}(k)$, one of the main tools which used is a result in homological algebra on the vanishing of Ext groups:
@@
Other key results applied in the proof of Theorem 1.1 are Lemmas 2.1 and 3.1. As an immediate application of this theorem, the desired splitting is deduced:
@@
We also show that the above exact sequence does not split for most incomplete local rings (see Example 3.19).
@@ 2. Finite case-the initial motivation
The above lemma, in particular, tells us that we have an isomorphism of Abelian groups $R^{\ast}/(1+I)\simeq(R/I)^{\ast}$.
It can be seen that the characteristic of any ring $R$ with only trivial idempotents is zero or a power of a prime number (local rings as well as rings with a unique minimal prime ideal are typical examples of such rings). Indeed, if $\operatorname{char}(R)=n$ is not in the above form, then using Chinese remainder theorem, the prime subring of $R$ (that is isomorphic to $\mathbb{Z}/n$) and so $R$ will have non-trivial idempotents which is a contradiction. Also recall that if $f:A\rightarrow B$ is a morphism of rings, then it can be seen that $\operatorname{char}(B)$ divides $\operatorname{char}(A)$. In particular, if we have morphisms of rings $A\rightarrow B$ and $B\rightarrow A$, then $\operatorname{char}(A)=\operatorname{char}(B)$. If we have a ring map $A\rightarrow B$ with $\operatorname{char}(B)=0$, then $\operatorname{char}(A)=0$. However, if $\operatorname{char}(A)=p^{n}$ for some prime number $p$, then $\operatorname{char}(B)=p^{d}$ with $d\leqslant n$. Keeping these in mind, we come to the following conclusion which is a well known result, its only innovation is probably our new proof:
@@
Let $(R,M)$ be a finite local ring of order $n$. We know that $R$ is of characteristic $p^{d}$ where $p$ is a prime number and $d\geqslant 1$. If $q$ is a prime factor of $n$ then by Cauchy theorem for finite groups (see e.g. [6, Corollary 5.4.]), the additive group of $R$ has an element $x\in R$ of order $q$. This means that $q\cdot x=0$. We also have $p^{d}\cdot x=0$. Then $q|p^{d}$ and so $q=p$. This shows that $n=p^{k}$ where $k\geqslant d$. Finally, using the natural ring map $R\rightarrow k$ where $k=R/M$ is the residue field of $R$, we conclude that $p=\operatorname{char}(k)$. ∎
@@
Let $G$ be a finite Abelian group of order $|G|=mn$ with $\gcd(m,n)=1$. If $H$ is a subgroup of $G$ of order $m$, then there exists a subgroup $K$ of $G$ of order $n$ such that $G=HK\simeq H\times K$ and the following exact sequence:
@@
This is an easy exercise in basic group theory. In fact, it can be seen that $H=\{x\in G:x^{m}=1\}$. Indeed, take $x\in G$ with $x^{m}=1$. If $x\notin H$ then $xH$ is a nontrivial element of the group $G/H$ and hence its order $d\geqslant 2$. But $|G/H|=n$ and so $d|n$. We also have $(xH)^{m}=H$ and so $d|m$. But this is in contradiction with $\gcd(m,n)=1$. Since $G$ is Abelian, $K:=\{x\in G:x^{n}=1\}$ is a subgroup of $G$. We also have $H\cap K=1$, because $\gcd(m,n)=1$. Thus $|HK|=(|H|\cdot|K|)/|H\cap K|=mn$. This shows that $G=HK$. Then the map $G=HK\rightarrow H\times K$ given by $g=hk\mapsto(h,k)$ is an isomorphism of groups, and its composition with the projective map $H\times K\rightarrow H$ is an splitting of the inclusion map $H\rightarrow G$. ∎
@@ 3. The general case
This section is the technical heart of this article. To prove the main results, we need the following lemmas:
@@
Let $(R,M,k)$ be a local ring. Then for each $n\geqslant 1$, the multiplicative group $A=(1+M^{n})/(1+M^{n+1})$ is canonically isomorphic to the additive group $M^{n}/M^{n+1}$, and we have the following short exact sequence of Abelian groups:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/M^{n+1})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/M^{n})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
In particular, the group $A$ is vanished by $\operatorname{char}(k)$.
@@
We first show that the map $A\rightarrow M^{n}/M^{n+1}$ given by $(1+x)(1+M^{n+1})\mapsto x+M^{n+1}$ is well-defined. If $(1+x)(1+M^{n+1})=(1+y)(1+M^{n+1})$ for some $x,y\in M^{n}$, then $(1+x)z\in 1+M^{n+1}$ where $z=(1+y)^{-1}$. We have $z=1-yz\in 1+M^{n}$. It follows that $(x-y)z\in M^{n+1}$. But $M^{n+1}$ is a primary ideal of $R$ and $z\notin\sqrt{M^{n+1}}=M$. Thus $x-y\in M^{n+1}$. This map is clearly bijective. If $x,y\in M^{n}$ then $(1+x)(1+y)(1+M^{n+1})=(1+x+y)(1+M^{n+1})$, because $xy\in M^{2n}\subseteq M^{n+1}$. This shows that the above map is also a group morphism. Hence, this map is an isomorphism of Abelian groups: $A\simeq M^{n}/M^{n+1}$. To see the exactness of the above sequence, it can be seen that the map $A\rightarrow 1+(M^{n}/M^{n+1})\subseteq(R/M^{n+1})^{\ast}$ given by $(1+x)(1+M^{n+1})\mapsto(1+x)+M^{n+1}$ is an isomorphism of groups, and then apply Lemma 2.1. The statement about the vanishing is clear, because $A\simeq M^{n}/M^{n+1}$ is a $k$-module, and it is obvious that if $N$ is a module over a ring $R$ then $N$ is vanished by $\operatorname{char}(R)$. ∎
@@
Let $\mathscr{C}$ be a category which admits (filtered) direct limits. Then for any object $B$ and for any direct system of objects $(A_{i})_{i\in I}$ in this category, it can be easily seen that the $\operatorname{Hom}_{\mathscr{C}}(A_{i},B)$ is an inverse system of sets over the same index set $I$ and the canonical map:
$\operatorname{Hom}_{\mathscr{C}}(\lim\limits_{\overrightarrow{i\in I}}A_{i},B)\rightarrow\lim\limits_{\overleftarrow{i\in I}}\operatorname{Hom}_{\mathscr{C}}(A_{i},B)$
given by $f\mapsto(f\pi_{i})_{i\in I}$ is an isomorphism (bijection) of sets where $\pi_{k}:A_{k}\rightarrow\lim\limits_{\overrightarrow{i\in I}}A_{i}$ is the canonical morphism. Similarly, let $\mathscr{C}$ be a category which admits inverse limits. Then for any object $A$ and for any inverse system of objects $(B_{i})_{i\in I}$ in this category, it can be easily seen that the $\operatorname{Hom}_{\mathscr{C}}(A,B_{i})$ is an inverse system of sets over the same index set $I$ and the canonical map:
@@
given by $f\mapsto(p_{i}f)_{i\in I}$ is an isomorphism (bijection) of sets where $p_{k}:\lim\limits_{\overleftarrow{i\in I}}B_{i}\rightarrow B_{k}$ is the projection morphism. In addition, if $\mathscr{C}$ is the category of Abelian groups (or the category of $R$-modules) then the above maps are indeed isomorphisms of Abelian groups ($R$-modules).
@@
can be fulfilled, i.e., one can find a unique ring map $S\rightarrow A$ that makes the new diagram commute, where $I$ is a nilpotent ideal of $A$ and $A\rightarrow A/I$ is the canonical ring map. To find such a ring map, we first extend from $A/I$ to $A/I^{2}$, then to $A/I^{4}$ and so on. More precisely, the following diagram is commutative:
where $A/I^{2}\rightarrow A/I$ is the canonical ring map. Since the ideal $I/I^{2}\subseteq A/I^{2}$ is of square-zero, there exists a unique ring map $S\rightarrow A/I^{2}$ making the new diagram commute. Then apply the same argument for the following commutative diagram:
@@
where $I$ is an ideal of $A$ of square-zero and $\pi$ is the canonical ring map. Then for each $i$, the following diagram is commutative:
where $\alpha_{i}:S_{i}\rightarrow S$ is the canonical ring map, and so there exists a unique ring map $g_{i}:S_{i}\rightarrow A$ making the new diagram commute. It can be easily seen that the maps $g_{i}$ are compatible with the transition maps of the above direct system. Then by the universal property of direct limits, there exists a unique ring map $g:S\rightarrow A$ such that $g_{i}=g\alpha_{i}$ for all $i$. We show that $g$ is the unique ring map making the above first diagram commute. We may assume that $S$ is a nonzero ring (if $S$ is a zero ring, then $A$ is also a zero ring and the assertion is clear in this case), and hence the index set of the above direct system is nonempty, so we may choose some index $i$. We have $f=\alpha_{i}f_{i}$ and so $gf=g\alpha_{i}f_{i}=g_{i}f_{i}=h$. To see $\pi g=h^{\prime}$, take $x\in S$, then there exists an index $i$ and some $s\in S_{i}$ such that $x=\alpha_{i}(s)$, then $(\pi g)(x)=\pi g\alpha_{i}(s)=\pi g_{i}(s)=h^{\prime}\alpha_{i}(s)=h^{\prime}(x)$. Hence, $g$ makes the above first diagram commute. To prove its uniqueness, suppose $g^{\prime}:S\rightarrow A$ is another ring map making the above first diagram commute. To see $g=g^{\prime}$, it suffices to show that $g_{i}=g^{\prime}\alpha_{i}$ for all $i$. We have $g^{\prime}\alpha_{i}f_{i}=g^{\prime}f=h$ and $\pi g^{\prime}\alpha_{i}=h^{\prime}\alpha_{i}$, and so by the uniqueness of $g_{i}$, we get that $g_{i}=g^{\prime}\alpha_{i}$ for all $i$. This completes the proof. ∎
@@
Also recall that the notion of separable extension can be also defined for arbitrary extension of fields (see e.g. [2, Tag 030O]). We mention this definition here for the convenience of the reader:
@@
An arbitrary extension of fields $K/k$ (not necessarily algebraic) is called a separable extension if for any subextension $k\subseteq K^{\prime}\subseteq K$ with $K^{\prime}$ a finitely generated field extension of $k$, the extension $k\subseteq K^{\prime}$ has a separating transcendence basis.
@@
It is well known that an arbitrary extension of fields is separable (in the above sense) if and only if this extension is formally smooth (see [2, Tags 031Y and 0320]). The following theorem improves this important result in the algebraic extension case:
Theorem 3.7.
For an algebraic extension of fields $K/k$ the following statements are equivalent:
$\mathbf{(i)}$ The extension $K/k$ is formally étale.
$\mathbf{(ii)}$ The extension $K/k$ is formally smooth.
$\mathbf{(iii)}$ The extension $K/k$ is a separable algebraic extension.
@@
[…]
@@
In relation with the above theorem, it is important to note that although any flat and separable morphism of commutative rings is formally étale [4, Corollary 4.7.3], the notion of separable morphism of rings in the sense of [4] only generalizes to “finite” separable extensions in fields (see [4, Corollary 4.5.4(1)]).
@@
In the characteristic zero, every algebraic extension of fields is separable. Then apply Theorem 3.7. ∎
To prove the first main theorem we also need to the following key result:

### v3

@@ Abstract.
Let $(R,M,k)$ be a complete local ring (not necessarily Noetherian). Then we prove that the natural surjective ring map $R\rightarrow k$ admits a splitting if and only if $\operatorname{char}(R)=\operatorname{char}(k)$. In addition, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then we prove that the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
We also show with an example that the above theorem is in its full generality. As an application of this theorem, we show that the following short exact sequence of Abelian groups:
@@
is always split. Next, we show with an example that the above exact sequence does not split for many incomplete local rings.
@@
complete local ring; the group of units; splitting; characteristic $p$ methods; (separating) transcendence basis; formally étale ring map; formally smooth ring map; Ext groups; pullbacks in the category of (commutative) rings
@@ 1. Introduction
In mathematics, it is common for a particular case of a hidden fact to appear first of all. Then, as the obstacles and coverings are gradually removed, the general form of that phenomenon becomes fully revealed. The same happened in this research work. More precisely, by elementary methods we first showed that for any “finite” local ring $(R,M,k)$ the following short exact sequence of Abelian groups:
@@
is split. Next, this observation naturally led us to the question of whether the same splitting holds for arbitrary (infinite) local rings.
Then, after exploring several stages of the level of generalization, first for Artinian local rings, then for local rings whose Jacobson radical are nilpotent, finally as one of the main results of this investigation, we succeeded to show that the same splitting holds for any complete local ring. However, the proof of this general case is technical, and uses the tools and methods of commutative algebra, the theory of fields, and homological algebra. In fact, we prove a more general result:
@@
Let $(R,M,k)$ be a complete local ring. Then the natural surjective ring map $R\rightarrow k$ admits a splitting if and only if $\operatorname{char}(R)=\operatorname{char}(k)$. Furthermore, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the natural surjective map between the groups of units $R^{\ast}\rightarrow k^{\ast}$ admits a splitting.
In proving this theorem, as is clear from its formulation, the characteristic of the residue field $k=R/M$ plays an important role and hence it is necessary to examine it carefully. We know that in any local ring $(R,M,k)$, if $\operatorname{char}(R)=\operatorname{char}(k)$ then there are two cases: $\operatorname{char}(k)=0$ or $\operatorname{char}(k)=p$ for some prime number $p$. However, if $\operatorname{char}(R)\neq\operatorname{char}(k)$, then the characteristic of $k$ is always a prime number $p$. To prove the above theorem, in the first case, i.e., when $\operatorname{char}(R)=\operatorname{char}(k)$, one of the main ingredients that applied is the following important result:
@@
Then the following result paves the way for proving Lemma 1.2, especially in the positive characteristic case.
Lemma 1.4.
@@
To prove Theorem 1.1, in the second case, i.e., when $\operatorname{char}(R)\neq\operatorname{char}(k)$, one of the main tools which used is a key result in homological algebra on the vanishing of Ext groups:
@@
Other results applied in the proof of Theorem 1.1 are Lemmas 2.1 and 3.1. As an immediate application of this theorem, the desired splitting is deduced:
@@
Next, we show that the above exact sequence does not split for most incomplete local rings (see Example 3.18).
We also show with an example that Theorem 1.1 is in its full generality, and it cannot be generalized further (see Example 3.25).
@@ 2. Finite case-the initial motivation
It can be seen that the characteristic of any ring $R$ with only trivial idempotents is zero or a power of a prime number (local rings as well as rings with a unique minimal prime ideal are typical examples of such rings). Indeed, if $n=\operatorname{char}(R)$ is not in the above form, then the prime subring of $R$ which is isomorphic to $\mathbb{Z}/n$ will have non-trivial idempotents (in fact, using Chinese remainder theorem, this ring has precisely $2^{d}$ idempotents where $d$ is the number of distinct primes that appear in the prime factorization of $n$). Then $R$ will also have non-trivial idempotents, which is a contradiction.
Also recall that if $f:A\rightarrow B$ is a morphism of rings, then it can be seen that $\operatorname{char}(B)$ divides $\operatorname{char}(A)$. In particular, if we have morphisms of rings $A\rightarrow B$ and $B\rightarrow A$, then $\operatorname{char}(A)=\operatorname{char}(B)$. If we have a ring map $A\rightarrow B$ with $\operatorname{char}(B)=0$, then $\operatorname{char}(A)=0$. However, if $\operatorname{char}(A)=p^{n}$ for some prime number $p$, then $\operatorname{char}(B)=p^{d}$ with $d\leqslant n$.
Keeping these in mind, we come to the following conclusion which is a well known result, its only innovation is probably our new proof:
@@
Let $(R,M,k)$ be a finite local ring of order $n$. We know that $R$ is of characteristic $p^{d}$ where $p$ is a prime number and $d\geqslant 1$. If $q$ is a prime factor of $n$ then by Cauchy theorem for finite groups (see e.g. [6, Corollary 5.4.]), the additive group of $R$ has an element $x\in R$ of order $q$. This means that $q\cdot x=0$. We also have $p^{d}\cdot x=0$. Then $q|p^{d}$ and so $q=p$. This shows that $n=p^{k}$ where $k\geqslant d$. Finally, using the natural ring map $R\rightarrow k$ we conclude that $p=\operatorname{char}(k)$. ∎
@@
Let $G$ be a finite Abelian group of order $|G|=mn$ with $\gcd(m,n)=1$. If $H$ is a subgroup of $G$ of order $m$, then there exists a subgroup $K$ of $G$ of order $n$ such that $G=HK\simeq H\times K$ and the following short exact sequence:
@@
This is an easy exercise in basic group theory. In fact, it can be seen that $H=\{x\in G:x^{m}=1\}$. Indeed, take $x\in G$ with $x^{m}=1$. If $x\notin H$ then $xH$ is a nontrivial element of the group $G/H$ and hence its order $d\geqslant 2$. But $|G/H|=n$ and so $d|n$ (because according to Lagrange’s theorem, the order of each member divides the order of the group). We also have $(xH)^{m}=H$ and so $d|m$. But this is in contradiction with $\gcd(m,n)=1$.
Since $G$ is Abelian, $K:=\{x\in G:x^{n}=1\}$ is a subgroup of $G$. We also have $H\cap K=1$, because $\gcd(m,n)=1$. Thus $|HK|=(|H|\cdot|K|)/|H\cap K|=mn$. This shows that $G=HK$. Then the map $G=HK\rightarrow H\times K$ given by $g=hk\mapsto(h,k)$ is an isomorphism of groups, and its composition with the projective map $H\times K\rightarrow H$ is an splitting of the inclusion map $H\rightarrow G$. ∎
@@
Recall that by a complete local ring we mean a local ring $(R,M)$ such that it is complete with respect to the $M$-adic topology (i.e., $R$ is Hausdorff and every Cauchy sequence is convergent in $R$), or equivalently, the natural ring map $R\rightarrow\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ given by $r\mapsto(r+M^{n})_{n\geqslant 1}$ is an isomorphism. Note that in this definition $R$ is not assumed to be Noetherian (contrary to the classical literature, see e.g. [1], where complete local rings were also assumed to be Noetherian).
Remark 2.5.
It can be easily seen that every Artinian local ring or more generally every local ring whose maximal ideal is nilpotent is a complete local ring. In particular, every finite local ring is complete. Let $M$ be a maximal ideal of a ring $R$. It is clear that the maximal ideal $M/M^{d}$ of $R/M^{d}$ is nilpotent and hence it is a complete local ring for all $d\geqslant 1$. It can be also easily seen that $S=\lim\limits_{\overleftarrow{n\geqslant 1}}R/M^{n}$ is a local ring with the maximal ideal $\mathfrak{m}:=\operatorname{Ker}(S\rightarrow R/M)$ and the residue field $R/M$. If moreover, $M$ is a finitely generated ideal of $R$, then $S$ is a complete local ring, $\mathfrak{m}=MS$ and we have a natural isomorphism of rings $S/\mathfrak{m}^{d}\simeq R/M^{d}$ for all $d\geqslant 1$ (see [2, 05GG]). However, in general, the local ring $S$ is not necessarily complete with respect to the $\mathfrak{m}$-adic topology (see e.g. [2, Tag 05JA]).
@@ 3. The general case
This section is the technical heart and culmination of this article. To prove the main results, we need the following lemmas:
@@
Let $I$ be an ideal of a ring $R$ which is contained in the Jacobson radical of $R$. Then for each $n\geqslant 1$, the multiplicative group $G=(1+I^{n})/(1+I^{n+1})$ is naturally isomorphic to the additive group $I^{n}/I^{n+1}$, and we have the following short exact sequence of Abelian groups:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${G\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/I^{n+1})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/I^{n})^{\ast}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
In particular, the Abelian group $G$ is vanished by $\operatorname{char}(R/I)$.
@@
We know that $1+I^{n}$ is a subgroup of $R^{\ast}$ (see Lemma 2.1). The map $1+I^{n}\rightarrow I^{n}/I^{n+1}$ given by $1+x\mapsto x+I^{n+1}$ is a morphism of groups, because if $x,y\in I^{n}$ then $xy\in I^{2n}\subseteq I^{n+1}$ and so $x+y+xy+I^{n+1}=x+y+I^{n+1}$. This map is clearly surjective and its kernel is equal to $1+I^{n+1}$. Hence, the above map induces an isomorphism of Abelian groups $G\simeq I^{n}/I^{n+1}$.
To see the exactness of the above sequence, it can be seen that the map $1+I^{n}\rightarrow 1+(I^{n}/I^{n+1})\subseteq(R/I^{n+1})^{\ast}$ given by $1+x\mapsto(1+x)+I^{n+1}$ is a surjective morphism of groups whose kernel is equal to $1+I^{n+1}$. Thus this map induces an isomorphism of Abelian groups $G\simeq 1+(I^{n}/I^{n+1})$. Then apply Lemma 2.1.
Using the above isomorphism of Abelian groups we can make $G$ into an $R/I$-module (and then this map becomes an isomorphism of $R/I$-modules). Now the statement about the vanishing is clear, because if $M$ is a module over a ring $R$ then $M$ is vanished by $\operatorname{char}(R)$. ∎
@@
Let $\mathscr{C}$ be a category which admits inverse limits. Then for any object $A$ and for any inverse system of objects $(B_{i})_{i\in I}$ in this category, it can be easily seen that the $\operatorname{Hom}_{\mathscr{C}}(A,B_{i})$ is an inverse system of sets over the same index set $I$ and the natural map:
@@
given by $f\mapsto(p_{i}f)_{i\in I}$ is an isomorphism (bijection) of sets where $p_{k}:\lim\limits_{\overleftarrow{i\in I}}B_{i}\rightarrow B_{k}$ is the projection morphism.
@@
can be fulfilled, i.e., one can find a unique ring map $S\rightarrow A$ that makes the new diagram commute, where $I$ is a nilpotent ideal of $A$ and $A\rightarrow A/I$ is the natural ring map. To find such a ring map, we first extend from $A/I$ to $A/I^{2}$, then to $A/I^{4}$ and so on. More precisely, the following diagram is commutative:
where $A/I^{2}\rightarrow A/I$ is the natural ring map. Since the ideal $I/I^{2}\subseteq A/I^{2}$ is of square-zero, there exists a unique ring map $S\rightarrow A/I^{2}$ making the new diagram commute. Then apply the same argument for the following commutative diagram:
@@
where $I$ is an ideal of $A$ of square-zero and $\pi$ is the natural ring map. Then for each $i$, the following diagram is commutative:
where $\alpha_{i}:S_{i}\rightarrow S$ is the natural ring map, and so there exists a unique ring map $g_{i}:S_{i}\rightarrow A$ making the new diagram commute. It can be easily seen that the maps $g_{i}$ are compatible with the transition maps of the above direct system. Then by the universal property of direct limits, there exists a unique ring map $g:S\rightarrow A$ such that $g_{i}=g\alpha_{i}$ for all $i$.
We show that $g$ is the unique ring map making the above first diagram commute. We may assume that $S$ is a nonzero ring (if $S$ is a zero ring, then $A$ is also a zero ring and the assertion is clear in this case), and hence the index set of the above direct system is nonempty, so we may choose some index $i$. We have $f=\alpha_{i}f_{i}$ and so $gf=g\alpha_{i}f_{i}=g_{i}f_{i}=h$. To see $\pi g=h^{\prime}$, take $x\in S$, then there exists an index $i$ and some $s\in S_{i}$ such that $x=\alpha_{i}(s)$, then $(\pi g)(x)=\pi g\alpha_{i}(s)=\pi g_{i}(s)=h^{\prime}\alpha_{i}(s)=h^{\prime}(x)$. Hence, $g$ makes the above first diagram commute. To prove its uniqueness, suppose $g^{\prime}:S\rightarrow A$ is another ring map making the above first diagram commute. To see $g=g^{\prime}$, it suffices to show that $g_{i}=g^{\prime}\alpha_{i}$ for all $i$. We have $g^{\prime}\alpha_{i}f_{i}=g^{\prime}f=h$ and $\pi g^{\prime}\alpha_{i}=h^{\prime}\alpha_{i}$, and so by the uniqueness of $g_{i}$, we get that $g_{i}=g^{\prime}\alpha_{i}$ for all $i$. This completes the proof. ∎
@@
Also recall that the notion of separable extension can be also defined for arbitrary extension of fields (see e.g. [2, Tag 030O]). For convenience, we mention this definition here:
@@
An arbitrary extension of fields $K/k$ (not necessarily algebraic) is called separable if for any subextension $k\subseteq K^{\prime}\subseteq K$ with $K^{\prime}$ a finitely generated field extension of $k$, the extension $k\subseteq K^{\prime}$ has a separating transcendence basis.
@@
Note that Definition 3.5 can in turn be generalized in the following way (although we do not need it in this article): a morphism of rings $k\rightarrow A$ with $k$ a field is called a separable morphism (in the sense of [8, §26]) if for any field extension $k^{\prime}/k$ the ring $k^{\prime}\otimes_{k}A$ is reduced. For field extensions, these two definitions are essentially the same. More precisely, an extension of fields $K/k$ is separable (in the sense of Definition 3.5) if and only if it is formally smooth (see [2, Tags 031Y and 0320]), or equivalently, for any field extension $k^{\prime}/k$ the ring $k^{\prime}\otimes_{k}K$ is reduced (see [8, Theorem 26.9]). In the algebraic extension case, we have the following result:
Lemma 3.7.
An algebraic extension of fields $K/k$ is separable if and only if it is formally étale.
@@
First assume that $K/k$ is a separable algebraic extension of fields. Then we can write $K$ as the filtered union (direct limit) of all finite separable subextensions of $k$, i.e., $K=\bigcup_{\alpha}K_{\alpha}=\lim\limits_{\overrightarrow{\alpha}}K_{\alpha}$ where each subfield $K_{\alpha}$ of $K$ is a finite separable extension of $k$ (note that the $K_{\alpha}$ are directed by inclusion inside $K$ and the transition maps are indeed the inclusion maps). It is well known that every finite separable extension of fields is formally étale (see [2, Tags 00U3 and 00UR] or [4, Corollary 4.7.3]). Thus for each $\alpha$, the extension $K_{\alpha}/k$ is formally étale. Then by Lemma 3.4, then original extension $K/k$ is formally étale as well. Conversely, if the extension $K/k$ is formally étale, then by [2, Tag 031Y], this extension is separable in the above sense (see Definition 3.5). Then by Remark 3.6, it is a separable algebraic extension. ∎
@@
In relation with the above theorem, note that although any flat and separable morphism of commutative rings is formally étale [4, Corollary 4.7.3], the notion of separable morphism of rings in the sense of [4] only generalizes to “finite” separable extensions in fields (see [4, Corollary 4.5.4(1)]).
@@
In the characteristic zero, every algebraic extension of fields is separable. Then apply Lemma 3.7. ∎
To prove the first main theorem we also need to the following result: